OpenStudy (anonymous):

Please help! Solve for x. lnx(x^2-3x-4)+lnx(x^2-9x+20)=0

OpenStudy (anonymous):

divide both sides by lnx?

OpenStudy (anonymous):

@Mertsj you up to helping me out? the other guy from earlier said he didnt know how to do it without a calculator

OpenStudy (anonymous):

why don't you divide both sides by lnx? unless you meant to write something different

OpenStudy (anonymous):

im not sure you can do that. (oh i wasnt talking about you btw)

OpenStudy (anonymous):

if what you meant is lnx times (x^2-3x-4) then you can factor it out

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

there are two ways you can approach this lnx(x^2-3x-4)+lnx(x^2-9x+20)=0 either divide everything by lnx (x^2-3x-4)+(x^2-9x+20)=0 or lnx(x^2-3x-4)=-lnx(x^2-9x+20) and raise to the e to get x(x^2-3x-4)=-x(x^2-9x+20)

OpenStudy (anonymous):

ok one second let me try the first way

OpenStudy (anonymous):

um nevermind idk how to do it -_-

OpenStudy (anonymous):

(x^2-3x-4)+(x^2-9x+20)=0 2x^2-12x+16=0 divide by 2 x^2-6x+8=0 factorise to get (x-4)(x-2)=0 x=2 or x=4

OpenStudy (anonymous):

ok that makes sense but the answer key has x=1 as an answer as well.

OpenStudy (anonymous):

in that case i think i'm misunderstanding the question. could you type it using the equation signs?

OpenStudy (anonymous):

\[\ln x(x^2-3x-4)+lnx(x^2-9x+20)=0\]

OpenStudy (anonymous):

2 and 4 are answers as well, but one is also

OpenStudy (anonymous):

ah of course it is! we have factored out lnx which is also a solution. silly me! lnx = 0 e ^ lnx = e ^ 0 x = 1

OpenStudy (anonymous):

ahh ok! thanks for your time!

OpenStudy (anonymous):

no problem ;)

OpenStudy (anonymous):

\[xy _{1}+xy _{1}y^2-3=5y _{1}+xy \] solve for \[y _{1}\]